# Pluricanonical Geometry of Varieties Isogenous to a Product: Chevalley-Weil Theory and Pluricanonical Decompositions of Abelian Covers

### v1

@@ 5. Pluricanonical maps of a VIP
The first is a necessary criterion, which uses the character $\chi_{m}$ of the $m$-canonical representation $\rho_{m}$. The second is a sufficient criterion which additionally requires the building data of the abelian cover $\pi$. When $Y=\mathbb{P}^{1}\times\dotso\times\mathbb{P}^{1}$, the second criterion becomes particularly easy to apply, especially for regular unmixed VIPs with abelian group.
5.1. Generalities
[…]
(13) $Z_{m}\cap V(x_{0},\dotso,x_{k})=\Phi_{m}^{Z}\left(V(s_{0},\dotso,s_{k})\right).$
Proposition 5.1.
Proof.
Note that $K_{X}^{\otimes m}$ being basepoint free amounts to saying that $V(s_{0},\dotso,s_{k})=\emptyset$. This is equivalent by (13) to $Z_{m}$ not meeting the centre of the projection pp. In other words, $\Phi_{m}^{X}$ is a morphism iff $p_{|Z_{m}}$ is a morphism. In this case $p_{|Z_{m}}$ is a finite morphism, therefore from diagram (12) it follows that $\Phi_{m}^{X}$ must be finite. ∎
Corollary 5.2.
Proposition 5.3.
  1. (1)
  2. (2)
    If one of the curves $C_{i}$ has genus $2$, then $\Phi_{2}^{X}$ is not birational.
Proof.
(1) Suppose w.l.o.g. that $C_{1}$ is hyperelliptic. Then the map $\Phi_{1}^{C_{1}}\colon C_{1}\to\mathbb{P}^{1}$ is a double cover, thus $Z_{1}\cong\mathbb{P}^{1}\times\overset{\sim}{C_{2}}\times\dotso\times\overset{\sim}{C_{n}}$. If $\Phi_{1}^{X}$ is birational, then from diagram (12) we get a dominant rational map
Corollary 5.4.
Proof.
Since $K^{2}_{S}=8\chi(\mathcal{O}_{S})\geq 8$, Theorem 5.2 of [B-H-P-V] (alternatively, see [Bomb73]) implies that $\Phi_{m}^{S}$ is a birational morphism onto a normal surface. Thus, the result follows from Proposition 5.1. ∎
Remark 5.5.
By Theorem 1.1 of [CCZ07] the $m$-canonical map of a minimal Gorenstein projective 3-fold of general type with canonical singularities is birational for $m\geq 5$. This holds in particular for $3$-folds isogenous to a product since they are smooth and their canonical bundle is ample.
Now we investigate the birationality of $\Phi_{4}^{X}$ for a 3-fold $X$ isogenous to a product. For this purpose, we recall that a $(1,2)$ surface $S$ is a nonsingular projective surface of general type whose minimal model $S^{\prime}$ has $K^{2}_{S^{\prime}}=1$ and $p_{g}(S^{\prime})=2$. By [Hor76II, Theorem 4.8] a $(1,2)$ surface is simply connected.
Theorem 5.6.
Proof.
Since $p_{g}(X)\geq 5$, [CH21, Theorem 1.1] implies that $\Phi_{4}^{X}$ is birational iff there is no smooth 3-fold $Y$ birational to $X$ admitting a fibration $Y\to B$, where $B$ is a smooth curve and the general fibre $S$ is a $(1,2)$-surface. By contradiction, assume that such a fibration $Y\to B$ exists. Denote by $S^{\prime}$ the birational image in $X$ of the general fibre $S\subset Y$ and by $\hat{S}$ a resolution of singularities of $S^{\prime}$. Notice the $\hat{S}$ is simply connected since it is birational to the $(1,2)$ surface $S$. Hence, the map $\hat{S}\to S^{\prime}\subset X$ lifts to the universal cover of $X$, which is the product $\Delta^{3}$ of three unit disks:
[…]
5.2. Representation-Theoretic Methods
Let $\pi\colon X\to Y$ be a finite abelian cover between complex manifolds with Galois group $G$. In this subsection, we recall some birationality criteria for the canonical map of $X$, established by Catanese in [Cat18]. They use the character of the canonical representation of $\pi$. We present them here in the more general setting of $m$-canonical maps, using the characters $\chi_{m}$ of the pluricanonical representations
@@ 5.2. Representation-Theoretic Methods
The group algebra
@@
$\rho_{reg}\colon G\to\operatorname{GL}(\mathbb{C}[G]),\qquad h\mapsto\ \big\{\sum_{g\in G}\lambda_{g}g\ \mapsto\ \sum_{g\in G}\lambda_{g}(hg)\big\}.$
Given a general point $y\in Y$, we will construct a certain homomorphism between the representations $\rho_{m}$ and $\rho_{\mathrm{reg}}$. It turns out that its dual is closely related to the restriction of the $m$-canonical map $\Phi_{m}$ of $X$ to the fibre $\pi^{-1}(y)$.
@@
is a homomorphism of representations, i.e., for all $h\in G$ it holds $\epsilon_{y}\circ\rho_{m}(h)=\rho_{reg}(h)\circ\epsilon_{y}$.
@@
$\sum_{g\in G}\operatorname{ev}_{(h^{-1}g)(x_{0})}(h^{\ast}\omega)\cdot(h^{-1}g)^{-1}=\sum_{g\in G}\operatorname{ev}_{g(x_{0})}(\omega)\cdot hg^{-1}=\rho_{reg}(h)(\epsilon_{y}(\omega)).$
@@
Using the identification $G\cong\pi^{-1}(y)$ from (14), we obtain the following commutative diagram:
@@
where $\epsilon_{y}^{\vee}$ denotes the dual of the map $\epsilon_{y}$. In other words, the induced rational map
@@
Let $y\in U\subset Y$ be a general point as above and denote by $V_{m}$ the image $\epsilon_{y}(H^{0}(X,K_{X}^{\otimes m})\subset\mathbb{C}[G]$. We observe that we have a diagram as follows
@@
Thus, instead of $\epsilon_{y}^{\vee}$ we can consider the map $\mathbb{C}[G]^{\vee}\twoheadrightarrow V_{m}^{\vee}$, which we still call $\epsilon_{y}^{\vee}$ by abuse of notation, and the projectivizationfoot
@@
for every irreducible constituent $\chi$ of $\chi_{m}$. In other words, $\mathbb{P}(\epsilon_{y}^{\vee})(h)=\mathbb{P}(\epsilon_{y}^{\vee})(k)$ if and only if $\chi(h^{-1}k)=\chi^{\prime}(h^{-1}k)$ for all $\chi,\chi^{\prime}$ irreducible constituents of $\chi_{m}$. Since $h,k\in G$ are arbitrary, we are done.
@@
Next we present the sufficient criterion for the birationality of $\Phi_{m}$.
Proposition 5.13 (cf. Proposition 3.1 of [Cat18]).
@@
separates two distinct points $y=(y_{1},\dotso,y_{n})$ and $y^{\prime}=(y_{1}^{\prime},\dotso,y_{n}^{\prime})$ if and only if there exists $j$ such that $y_{j}\neq y_{j}^{\prime}$ and $r_{j}\geq 1$. Hence, if for every $j$ there is an irreducible constituent $\chi$ of $\chi_{m}$ such that $r_{m,\chi}^{j}\geq 1$, then every two points $y,y^{\prime}\in Y$ are separated, and this clearly implies condition $(b)$ of Proposition 5.13.
@@ 6. Examples and Computational Results
6.1. Example of a 3-Fold with Birational Canonical Map
@@
We begin with $C_{1}$. Applying the Chevalley-Weil formula [CheWeil34], we get the character of the canonical representation $H^{0}(C_{1},K_{C_{1}})$; since $G$ is abelian, we get right away the canonical representation itself:
@@
$e_{1}\mapsto\begin{pmatrix}-1&&&&\\ &-1&&&\\ &&1&&\\ &&&-1&\\ &&&&-1\\ \end{pmatrix},\quad e_{2}\mapsto\begin{pmatrix}-1&&&&\\ &-1&&&\\ &&-1&&\\ &&&1&\\ &&&&-1\\ \end{pmatrix},\quad e_{3}\mapsto\begin{pmatrix}-1&&&&\\ &-1&&&\\ &&-1&&\\ &&&-1&\\ &&&&1\\ \end{pmatrix}$.
@@
where $\mathcal{I}_{2}\subset\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4}]_{2}$ denotes the space of quadratic forms which vanish along $C_{1}$. Hence, the character $\chi_{\mathcal{I}_{2}}$ of $\mathcal{I}_{2}$ is the difference between the character of $\operatorname{Sym}^{2}(H^{0}(C_{1},K_{C_{1}}))$ and that of $H^{0}(C_{1},K_{C_{1}}^{\otimes 2})$, which we can compute by the Chevalley-Weil formula for pluricanonical representation, i.e., Theorem 3.4. It turns out that $\chi_{\mathcal{I}_{2}}=3\chi_{triv}$. In other words, the action of $G$ on $\mathcal{I}_{2}$ is trivial and $\dim_{\mathbb{C}}\mathcal{I}_{2}=3$. The space of $G$-invariant quadratic forms is
@@
(15) $\begin{cases}x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+ax_{0}x_{1}=0\\ x_{0}^{2}+x_{1}^{2}+x_{3}^{2}+bx_{0}x_{1}=0\\ x_{0}^{2}+x_{1}^{2}+x_{4}^{2}+cx_{0}x_{1}=0\end{cases}$
where $c\in\mathbb{C}\setminus\{\pm 2\}, aba\neq b,$ $a\neq c$, $b\neq c$. Indeed, the algebraic set $C_{a,b,c}\subset\mathbb{P}^{4}$ is a canonically embedded smooth projective curve of genus 5. Furthermore, the projection map
@@
From what we have said it follows that
@@
(16) $J:=\begin{pmatrix}z_{4}&0&x_{0}\\ y_{4}&x_{0}&0\\ \frac{\partial x_{1}}{\partial x_{0}}z_{4}&0&x_{1}\\ \frac{\partial x_{1}}{\partial x_{0}}y_{4}&x_{1}&0\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{0}&x_{2}z_{0}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{0}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{1}&x_{2}z_{1}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{1}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}z_{2}&0&x_{2}\frac{\partial z_{2}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{4}&y_{4}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{0}z_{2}&x_{4}z_{2}\frac{\partial y_{0}}{\partial y_{4}}&x_{4}y_{0}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{1}z_{2}&x_{4}z_{2}\frac{\partial y_{1}}{\partial y_{4}}&x_{4}y_{1}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{2}&x_{4}\frac{\partial y_{2}}{\partial y_{4}}&0\\ \end{pmatrix}$
@@ 6.2. Further Computational Results
Table 1. The 347 threefolds with $\chi=-1$ and $|G|\leq 16$
@@ 47 $\mathbb{Z}_{2}^{4}$ 2 2 4 $2^{5}$ $2^{6}$ $2^{6}$ 5 3 0 9 15 0 0 4 0 4 0 4 0
    with general fibre $C_{i}$ are not genus–$2$ fibrations. In fact, such a threefold does not admit any genus–$2$ fibration at all. Indeed, since $h^{2,0}(X)=h^{1,0}(X)=0$, we have
    $H^{1}(X,\mathcal{O}_{X}^{\ast})\simeq H^{2}(X,\mathbb{Z}).$
    Moreover, as $h^{1,1}(X)=3$, the threefold $X$ has Picard rank equal to 3. The fibres $F_{i}\simeq C_{j}\times C_{k}$ of the fibrations
    $\varphi_{i}\colon X\to C_{i}/G_{i}$
    generate $N^{1}(X)_{\mathbb{R}}$. They even generate the nef cone, namely
@@
    We conclude that $f_{1}$, $f_{2}$ and $f_{3}$ are the only fibrations of $X$ over a normal surface, since such fibrations correspond to the two-dimensional faces of the nef cone. In particular, $X$ does not admit any genus–$2$ fibration. Thus, this phenomenon closely resembles the so-called non-standard case for surfaces, see [BCP06, Section 2]) for an overview on the topic. Note that rows No. 40 and 44 provide exotic examples as well.
  3. (3)
    Finally, we observe that $K_{X}^{\otimes 2}$ is always base-point free and pose the question of whether this property holds in general for regular unmixed threefolds isogenous to a product. Furthermore, our birationality criteria are always satisfied for the 3-canonical map, even though this information is not included in the table. This motivates the question of whether $K_{X}^{\otimes 3}$ defines a birational map for all regular unmixed threefolds of this type. Recall that for a surface $S$ isogenous to a product $K_{S}^{\otimes 2}$ is base-point free by Reider’s theorem [Reider88] since $K^{2}_{S}\geq 8$; moreover, $K_{S}^{\otimes 3}$ defines an embedding (cf. Corollary 5.4).

### v2

@@ 4. Pluricanonical Maps of an Abelian Cover
The first one is a necessary criterion, which uses the character $\chi_{m}$ of the $m$-canonical representation
@@
The second one is a sufficient criterion which additionally requires the building data of the abelian cover $\pi$. When $Y=\mathbb{P}^{1}\times\dotso\times\mathbb{P}^{1}$, the second criterion becomes particularly easy to apply, especially for regular unmixed VIPs with abelian Galois group.
Let $\pi\colon X\to Y$ be a finite abelian cover between complex manifolds with Galois group $G$. The group algebra
@@
$\rho_{\mathrm{reg}}\colon G\to\operatorname{GL}(\mathbb{C}[G]),\qquad h\mapsto\ \big\{\sum_{g\in G}\lambda_{g}g\ \mapsto\ \sum_{g\in G}\lambda_{g}(hg)\big\}.$
Given a general point $y\in Y$, we will construct a certain homomorphism between the representations $\rho_{m}$ and $\rho_{\mathrm{reg}}$. It turns out that its dual map is closely related to the restriction of the $m$-canonical map $\Phi_{m}$ of $X$ to the fibre $\pi^{-1}(y)$.
@@
is a homomorphism of representations, i.e., for all $h\in G$ it holds $\epsilon_{y}\circ\rho_{m}(h)=\rho_{\mathrm{reg}}(h)\circ\epsilon_{y}$.
@@
$\sum_{g\in G}\operatorname{ev}_{(h^{-1}g)(x_{0})}(h^{\ast}\omega)\cdot(h^{-1}g)^{-1}=\sum_{g\in G}\operatorname{ev}_{g(x_{0})}(\omega)\cdot hg^{-1}=\rho_{\mathrm{reg}}(h)(\epsilon_{y}(\omega)).$
@@
Using the identification $G\cong\pi^{-1}(y)$ from (4.1), we obtain the following commutative diagram:
@@
where $\epsilon_{y}^{\vee}$ denotes the dual map of $\epsilon_{y}$. In other words, the induced rational map
@@
Let $y\in U\subset Y$ be a general point as above and denote by $V_{m}$ the image $\epsilon_{y}(H^{0}(X,K_{X}^{\otimes m}))\subset\mathbb{C}[G]$. We observe that we have a diagram as follows
@@
Thus, instead of $\epsilon_{y}^{\vee}$ we can consider the map $\mathbb{C}[G]^{\vee}\twoheadrightarrow V_{m}^{\vee}$, which we still call $\epsilon_{y}^{\vee}$ by abuse of notation, and the projectivization
@@
for every irreducible constituent $\chi$ of $\chi_{m}$. In other words, $\mathbb{P}(\epsilon_{y}^{\vee})(h)=\mathbb{P}(\epsilon_{y}^{\vee})(k)$ if and only if $\chi(h^{-1}k)=\chi^{\prime}(h^{-1}k)$ for all $\chi,\chi^{\prime}$ irreducible constituents of $\chi_{m}$. Since $h,k\in G$ are arbitrary, we are done. ∎
@@
Next we present the sufficient criterion for the birationality of $\Phi_{m}$ announced at the beginning of this section.
Proposition 4.7 (cf. [Cat18, Proposition 3.1]).
@@
separates two distinct points $y=(y_{1},\dotso,y_{n})$ and $y^{\prime}=(y_{1}^{\prime},\dotso,y_{n}^{\prime})$ if and only if there exists $j$ such that $y_{j}\neq y_{j}^{\prime}$ and $r_{j}\geq 1$. Hence, if for every $j$ there is an irreducible constituent $\chi$ of $\chi_{m}$ such that $r_{m,\chi}^{j}\geq 1$, then every two points $y,y^{\prime}\in Y$ are separated, and this clearly implies condition $(b)$ of Proposition 4.7. ∎
@@ 5. Examples and Computational Results
5.1. Example of a Threefold with Birational Canonical Map
@@
We begin with $C_{1}$. Applying the Chevalley-Weil formula [CheWeil34, Weil35], we get the character of the canonical representation $H^{0}(C_{1},K_{C_{1}})$; since $G$ is abelian, we get right away the canonical representation itself:
@@
$e_{1}\mapsto\operatorname{diag}(-1,-1,1,-1,-1),\qquad e_{2}\mapsto\operatorname{diag}(-1,-1,-1,1,-1),\qquad e_{3}\mapsto\operatorname{diag}(-1,-1,-1,-1,1).$
@@
where $\mathcal{I}_{2}\subset\mathbb{C}[x_{0},x_{1},x_{2},x_{3},x_{4}]_{2}$ denotes the space of quadratic forms which vanish along $C_{1}$. Hence, the character $\chi_{\mathcal{I}_{2}}$ of $\mathcal{I}_{2}$ is the difference between the character of $\operatorname{Sym}^{2}(H^{0}(C_{1},K_{C_{1}}))$ and that of $H^{0}(C_{1},K_{C_{1}}^{\otimes 2})$, which we can compute by the Chevalley-Weil formula for pluricanonical representations, i.e., Theorem 2.4. It turns out that $\chi_{\mathcal{I}_{2}}=3\chi_{triv}$. In other words, the action of $G$ on $\mathcal{I}_{2}$ is trivial and $\dim_{\mathbb{C}}\mathcal{I}_{2}=3$. The space of $G$-invariant quadratic forms is
@@
(5.1) $\left\{\begin{array}[]{l}x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+ax_{0}x_{1}=0\\ x_{0}^{2}+x_{1}^{2}+x_{3}^{2}+bx_{0}x_{1}=0\\ x_{0}^{2}+x_{1}^{2}+x_{4}^{2}+cx_{0}x_{1}=0\end{array}\right$.
where $c\in\mathbb{C}\setminus\{\pm 2\}, aba\neq b,$ $a\neq c$, $b\neq c$. Indeed, the algebraic set $C_{a,b,c}\subset\mathbb{P}^{4}$ is a canonically embedded smooth projective curve of genus 5.
Furthermore, the projection map
@@
From what we have said, it follows that
@@
(5.2) $J:=\begin{pmatrix}z_{4}&0&x_{0}\\ y_{4}&x_{0}&0\\ \frac{\partial x_{1}}{\partial x_{0}}z_{4}&0&x_{1}\\ \frac{\partial x_{1}}{\partial x_{0}}y_{4}&x_{1}&0\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{0}&x_{2}z_{0}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{0}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{1}&x_{2}z_{1}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{1}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}z_{2}&0&x_{2}\frac{\partial z_{2}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{4}&y_{4}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{0}z_{2}&x_{4}z_{2}\frac{\partial y_{0}}{\partial y_{4}}&x_{4}y_{0}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{1}z_{2}&x_{4}z_{2}\frac{\partial y_{1}}{\partial y_{4}}&x_{4}y_{1}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{2}&x_{4}\frac{\partial y_{2}}{\partial y_{4}}&0\\ \end{pmatrix}$
@@ 5.2. Further Computational Results
Table 1. The 347 families of threefolds with $\chi=-1$ and $|G|\leq 16$
@@ 47 $\mathbb{Z}_{2}^{4}$ 2 2 4 $2^{5}$ $2^{6}$ $2^{6}$ 5 3 0 9 15 0 0 4 0 4 0 4 0
    with general fibre $C_{i}$ are not genus-$2$ fibrations. Furthermore, Proposition 5.4 shows that $X$ does not carry any other fibration onto a surface since $h^{2,0}(X)=h^{1,0}(X)=0$ and $h^{1,1}(X)=3$. Thus, this phenomenon closely resembles the so-called non-standard case for surfaces, see [BCP06, Section 2] for an overview on the topic. Note that rows No. 40 and 44 provide exotic examples as well.
  3. (3)
    Finally, we observe that $K_{X}^{\otimes 2}$ is always base-point free and pose the question of whether this property holds in general for regular unmixed threefolds isogenous to a product. Furthermore, our birationality criteria are always satisfied for the 3-canonical map, even though this information is not included in the table. This motivates the question of whether $K_{X}^{\otimes 3}$ defines a birational map for all regular unmixed threefolds of this type. Recall that for a surface $S$ isogenous to a product $K_{S}^{\otimes 2}$ is base-point free by Reider’s theorem [Reider88] since $K^{2}_{S}\geq 8$; moreover, $K_{S}^{\otimes 3}$ defines an embedding (cf. Proposition 1.10).
Proposition 5.4.
Let $X=(C_{1}\times C_{2}\times C_{3})/G$ be a threefold isogenous to a product of unmixed type such that
$h^{2,0}(X)=h^{1,0}(X)=0\quad\makebox{and}\quad h^{1,1}(X)=3.$
Then, up to isomorphism, the only fibrations of $X$ onto surfaces or curves are
$f_{i}\colon X\to\prod_{j\neq i}C_{j}/G\cong\mathbb{P}^{1}\times\mathbb{P}^{1},\quad\pi_{i}\colon X\to C_{i}/G\cong\mathbb{P}^{1},\quad\makebox{$i=1,2,3.$}$
Moreover, the $f_{i}$’s correspond to the two-dimensional faces of $\mathrm{Nef}(X)$, while the $\pi_{i}^{\prime}s$ correspond to its one-dimensional faces, according to the Picard rank of the target.
Proof.
By using the exponential sequence, the conditions $h^{2,0}(X)=h^{1,0}(X)=0$ imply that
$H^{1}(X,\mathcal{O}_{X}^{\ast})\cong H^{2}(X,\mathbb{Z}).$
Moreover, since $h^{1,1}(X)=3$, the threefold $X$ has Picard rank $\rho(X)=3$. Hence, the general fibres $F_{i}\cong C_{j}\times C_{k}$ of the fibrations $\pi_{i}\colon X\to C_{i}/G$ generate $N^{1}(X)_{\mathbb{R}}$.
Claim 1.
The general fibres $F_{i}$ of the fibrations $\pi_{i}\colon X\to C_{i}/G_{i}$ generate the nef cone
@@
In particular, every $\mathbb{Q}$-divisor in $\operatorname{Nef}(X)$ is semiample.
Proof of Claim 1.
Let $D=a_{1}F_{1}+a_{2}F_{2}+a_{3}F_{3}$ be a nef divisor and let $C_{ij}:=F_{i}\cdot F_{j}$. We have
$0\leq D\cdot C_{ij}=a_{k}(F_{k}\cdot C_{ij})=a_{k}(F_{i}\cdot F_{j}\cdot F_{k}).$
Since $F_{i}\cdot F_{j}\cdot F_{k}>0$, we conclude that $a_{k}\geq 0$. This shows that $\operatorname{Nef}(X)$ is the cone spanned by the fibers $F_{i}$’s. The last part of the thesis follows immediately since the fibers $F_{i}$’s are semiample divisors. ∎
Now let $f\colon X\to Y$ be a fibration of $X$, i.e., a surjective morphism onto a normal projective variety $Y$ such that $f_{\ast}\mathcal{O}_{X}=\mathcal{O}_{Y}$ and assume $\dim Y\in\{1,2\}$. The condition $f_{\ast}\mathcal{O}_{X}=\mathcal{O}_{Y}$ implies that $f$ is induced by a linear system $|D|$, where $D:=f^{\ast}A$ and $A$ is very ample. Since $\dim Y>0$, the morphism $f$ is indeed the Iitaka fibration associated with $D$. Since $f\colon X\to Y$ is not an isomorphism, we have that $D\in\partial\mathrm{Nef}(X)$, i.e., $D$ belongs to one of the six extremal faces of the first octant of $\mathbb{R}^{3}\cong N^{1}(X)_{\mathbb{R}}$. Let $\mathcal{F}$ be the extremal face which contains $D$ in its relative interior, namely $D\in\overset{\circ}{\mathcal{F}}$.
Claim 2.
Let $C$ be a curve. If $D^{\prime}\in\overset{\circ}{\mathcal{F}}$, then $D^{\prime}\cdot C=0$ if and only if $D\cdot C=0.$
Proof of Claim 2.
Assume $D\cdot C=0$. Since $D\in\overset{\circ}{\mathcal{F}}$, there exists $G\in\overset{\circ}{\mathcal{F}}$ such that
$D\in(G,D^{\prime}):=\{tG+(1-t)D^{\prime}\mid t\in(0,1)\}.$
This yields
$0=D\cdot C=t_{0}\ \underset{\geq 0}{\underbrace{(G\cdot C)}}+(1-t_{0})\ \underset{\geq 0}{\underbrace{(D^{\prime}\cdot C)}},\quad 0<t_{0}<1,$
whence $D^{\prime}\cdot C=G\cdot C=0$. By symmetry, the claim follows. ∎
Note that Claim 1 implies that every $\mathbb{Q}$-divisor $D^{\prime}\in\operatorname{Nef}(X)\setminus\{0\}$ has an associated Iitaka fibration $f^{\prime}\colon X\to Y^{\prime}$. This is indeed the map induced by a linear system $|mD^{\prime}|$ for a suitable $m\in\mathbb{N}$. Assume that $D^{\prime}\in\overset{\circ}{\mathcal{F}}$. Claim 2 implies by Rigidity Lemma [Debarre, Lemma 1.15] that $f^{\prime}$ is isomorphic to $f\colon X\to Y$, i.e., there exists an isomorphism $g\colon Y\to Y^{\prime}$ such that $f^{\prime}=g\circ f$.
Claim 3.
$\mathcal{F}=f^{\ast}\operatorname{Nef}(Y).$
Proof of Claim 3.
Since $\operatorname{Nef}(X)\subseteq\mathbb{R}^{3}$ is the first octant, it suffices to consider only $\mathbb{Q}$-divisors.
Let $G\in\mathcal{F}$ be a $\mathbb{Q}$-divisor and let $f_{G}:X\to Z$ be its associated Iitaka fibration. Note that there exist $m\in\mathbb{N}$ and a very ample divisor $B$ on $Z$ such that $mG=f_{G}^{\ast}B$ is an integral divisor and $f_{G}$ is induced by the linear system $|mG|$. Now let $C$ be a curve. By applying the same argument given in the proof of Claim 2, we have that $D\cdot C=0$ implies $G\cdot C=0$. Hence, by Rigidity Lemma [Debarre, Lemma 1.15]) there is a morphism $h:Y\to Z$ such that $f_{G}=h\circ f$. This implies that
$mG=f^{\ast}_{G}B=f^{\ast}(h^{\ast}B)\in f^{\ast}\operatorname{Nef}(Y),$
whence $G\in f^{\ast}\operatorname{Nef}(Y)$.
Conversely, let $G\in f^{\ast}\operatorname{Nef}(Y)$ be a $\mathbb{Q}$-divisor. By projection formula, $G\cdot C=0$ for each curve $C$ contracted by $f$. Suppose by contradiction that $G\notin\mathcal{F}$. Since $f\colon X\to Y$ is not an isomorphism, it contracts some curves, hence $G\notin\operatorname{Amp}(X)$. This implies that there exists an extremal face $\mathcal{F}^{\prime}\neq\mathcal{F}$ such that $G\in\mathcal{F}^{\prime}$. Thus, there exists $H\in\operatorname{Amp}(X)$ such that $H\in(D,G)$. By the same argument given in the proof of Claim 2 we conclude that $H\cdot C=0$ for each curve $C$ contracted by $f$, a contradiction. Therefore, $G\in\mathcal{F}$. ∎
Since the pullback $f^{\ast}\colon N^{1}(Y)_{\mathbb{R}}\to N^{1}(X)_{\mathbb{R}}$ is injective and $\mathrm{Nef}(Y)\subseteq N^{1}(Y)_{\mathbb{R}}$ is a full-dimensional closed convex cone, Claim 3 implies that $\dim\mathcal{F}=\rho(Y).$
Finally, since the six fibrations $f_{i}$’s and $\pi_{i}$’s belong to the relative interior of pairwise different faces of the first octant $\operatorname{Nef}(X)\subseteq\mathbb{R}^{3}$, the given fibration $f:X\to Y$ coincides with one of those. ∎